OpenStudy (anonymous):

how do i solve for inverse tangent ..im working on tan theta = 14/11. im looking for theta

OpenStudy (hauntedwoodsgal):

I can help

OpenStudy (anonymous):

i was needing to do it without a calculator ..is it possible

OpenStudy (hauntedwoodsgal):

yes it is possible just put the question on the paper and do everything else that you could do on the calculator

OpenStudy (hauntedwoodsgal):

A Science calculator would really help

OpenStudy (hauntedwoodsgal):

besides the answer is simple

OpenStudy (anonymous):

okay so what if the full question is find \[\sin \theta\] and \[ \cos \theta \] as exact fractions given that \[0\le \theta \le \pi/2\] and \[\tan \theta= 14/11\]

OpenStudy (hauntedwoodsgal):

Ok do you know the answer??

OpenStudy (hauntedwoodsgal):

Oh and Welcome to Open Study it's a wonderful site :D we help alot

OpenStudy (anonymous):

so the answer for the inverse tan is 0.904

OpenStudy (hauntedwoodsgal):

I know the Answer to this: how do i solve for inverse tangent ..im working on tan theta = 14/11. im looking for theta

OpenStudy (hauntedwoodsgal):

No wait im wrong

OpenStudy (anonymous):

howd you do that ..that is what i dont get

OpenStudy (hauntedwoodsgal):

I have had this problem in a class before

OpenStudy (hauntedwoodsgal):

If you do it on a calculator it would say

OpenStudy (hauntedwoodsgal):

1.272727273

OpenStudy (anonymous):

ok...h0wd you get 4

OpenStudy (hauntedwoodsgal):

Oh dear hold on I have to draw it out.

OpenStudy (hauntedwoodsgal):



OpenStudy (hauntedwoodsgal):

Numerator: We call the top number the Numerator, it is the number at the top of the fraction you have. Denominator: We call the bottom number the Denominator, it is the whole number that is at the bottom, it is the number that is divided into.

OpenStudy (hauntedwoodsgal):

Step 1: Find a number you can multiply by the bottom of the fraction to make it 10, or 100, or 1000, or any 1 followed by 0s. Step 2: Multiply both the top and bottom by that number you choose. Step 3. Then write down just the top number, putting the decimal point in the correct spot, that is one space from the right hand side for every zero in the bottom number. a/b=c In the above expressions, a is called the dividend, b is called the divisor, and c is called the quotient; in the expression a/b, a is also called the numerator and b is also called the denominator This fraction calculator can also be used as fraction to percent, shopping discounts, coupon off, body fat, gross profit, weight loss, love, tax, population increase & decrease, sales profit. Once you know the values, to determine the % is easy .

OpenStudy (hauntedwoodsgal):

Ok you got it now???